Consider the following scenario:

Cute Camel Woodcraft Company’s income statement reports data for its first year of operation. The firm’s CEO would like sales to increase by 25% next year.

1. Cute Camel is able to achieve this level of increased sales, but its interest costs increase from 10% to 15% of earnings before interest and taxes (EBIT).
2. The company’s operating costs (excluding depreciation and amortization) remain at 70.00% of net sales, and its depreciation and amortization expenses remain constant from year to year.
3. The company’s tax rate remains constant at 40% of its pre-tax income or earnings before taxes (EBT).
4. In Year 2, Cute Camel expects to pay $300,000 and $2,306,475 of preferred and common stock dividends, respectively.

Complete the Year 2 income statement data for Cute Camel, then answer the questions that follow. Round each dollar value to the nearest whole dollar.

Cute Camel Woodcraft Company

Income Statement for Year Ending December 31

Year 1 Year 2 (Forecasted)
Net sales $30,000,000
Less: Operating costs, except depreciation and amortization 21,000,000
Less: Depreciation and amortization expenses 1,200,000 1,200,000
Operating income (or EBIT) $7,800,000
Less: Interest expense 780,000
Pre-tax income (or EBT) $7,020,000
Less: Taxes (40%) 2,808,000
Earnings after taxes $4,212,000
Less: Preferred stock dividends 300,000
Earnings available to common shareholders $3,912,000
Less: Common stock dividends 1,895,400
Contribution to retained earnings $1,605,525 $2,519,025• In Year 2, if Cute Camel has 25,000 shares of preferred stock issued and outstanding, then each preferred share should expe
Answer #1

Answer (1):

Completed the Year 2 income statement data for Cute Camel is as follows:

Cute Camel Woodcraft Company Income Statement for Years Ending December 31 Year 1 Net sales $30,000,000 Less: Operating costs

Above excel with 'show formula' is as follows:

Cute Camel Woodcraft Company Income Statement for Years Ending December 31 Year 1 Year 2 (Forecasted) | 5 Net sales 30000000

Answer (2):

(a) In year 2, if Cute Camel has 25,000 shares of preferred stock issued and outstanding , then each preferred share should expect to receive $12 in annual dividends.

(b) If cute Camel has 200,000 shares of common stock issued and oustanding, then the firm's earnings per share (EPS) is expected to change from $9.48 in Year 1 to $11.53 in Year 2.

(c) Cute Camel's before interest, taxes, depreciation and amortization (EBITDA) value changed from $9.00M in Year 1 to $11.25M in Year 2

(d). It is INCORRECT to say that Cute Camel's net inflows and outflows of cash at the end of years 1 & 2 are equal to the company's annual contribution to retained earnings, $1,605,525 and $2,519,025 respectively. This is because ALL BUT ONE of the item reported in the income statement involve payments & receipts of cash

Working:

(a) Annual dividend per preference share = 300000 / 25000 = $12

(b)

EPS year 1 = 1,895,400 / 200000 = $9.48

EPS year 2 = $2,306,475 / 200000 = $11.53

(c)

EBITDA Year 1 = EBIT + Depreciation and amortization = $7,800,000 + 1,200,000 = $9,000,000 = $9.00 million

EBITDA Year 2 = EBIT + Depreciation and amortization = $10,050,000 + 1,200,000 = $11,250,000 = $11.25 million